Comparative balance sheets for Softech Canvas Goods for 2018 and 2017 are shown below. Softech pays no dividends, and instead reinvests all earnings for future growth.

Comparative Balance Sheets
($ in 000s)
December 31
2018 2017
Assets:
Cash $ 107 $ 78
Accounts receivable 195 215
Short-term investments 88 78
Inventory 390 254
Property, plant, and equipment (net) 980 835
$ 1,760 $ 1,460
Liabilities and Shareholders’ Equity:
Current liabilities $ 430 $ 305
Bonds payable 255 255
Paid-in capital 685 685
Retained earnings 390 215
$ 1,760 $ 1,460


Required:
1. Determine the return on shareholders’ equity for 2018.
Answer #1

1. Return on shareholder equity = Net Income/Average Shareholder Equity

Net Income for 2018 = $390 - 215 = $175

Average Sharehoder's Equity =  (Opening+closing equity) / 2

= (685+390+685+215) / 2 = $987.50

Return on shareholder equity = 175 / 987.50 = 17.72%